∠x and ∠y are complementary. If m∠x = a + 3 and m∠y = 2a + 6, what is the measure of ∠x?
Answer by nerdybill(7384)

The sum of "complementary" angles is 90 degrees.
.
Therefore,
(a+3) + (2a+6) = 90
solving for 'a':
a+3 + 2a+6 = 90
3a+9 = 90
3a = 81
a = 27
.
since measure of angle 'a' is a+3:
a+3 = 27+3 = 30 degrees